OpenStudy (tester97):

What is the 50th term of the sequence that begins −4, 2, 8, 14…? How would i do this @ankit042

OpenStudy (anonymous):

As you may notice the difference between -4 and 2 is 6 and the difference between 2 and 8 is 6 and so on so this is an arithmetic sequence

OpenStudy (anonymous):

to find the nth term of an arithmetic sequence you would use the formula \[a_n=a_1+(n-1)d\] where n is the nth number you are finding and d is the difference

OpenStudy (ankit042):

a1 = -4 d = 6 nth term = a+(n-1)d

OpenStudy (nirmalnema):

290

OpenStudy (anonymous):

so now all you have to do is plug what you know and solve for what you don't\[a_n=-4+(50-1)6\]

OpenStudy (tester97):

\[49\times6+(-4)=290?\]

OpenStudy (nirmalnema):

yup

OpenStudy (tester97):

thanks guys